# Accurate Segmentation of Optic Disc and Cup from Multiple Pseudo-labels by Noise-Aware Learning

### 30 Nov 2023 v1

@@ II-C Final Loss
The total loss function follows the pattern of semi-supervised learning including supervised loss and unsupervised loss. In our approach i.e. clean pixels loss $\mathcal{L}_{c}$ and noisy pixels loss $\mathcal{L}_{n}$. The total loss is expressed as:
$\mathcal{L}_{total}=\alpha\mathcal{L}_{c}+\lambda\mathcal{L}_{n}.$ (7)
Empirically, we set $\alpha=1$ and $\lambda$ is a ramp-up trade-off weight commonly scheduled by the time-dependent Gaussian function [27] $\lambda(t)=w_{max}\cdot e^{(-5(1-\frac{t}{t_{max}})^{2})}$, where $w_{max}$ is the maximum weight commonly set as 0.1 [26] and $t_{max}$ is the maximum training iteration. Such a $\lambda$ weight representation avoids being dominated by misleading targets when starting online training.
@@ III-A Dataset and Pre-processing
The dataset is a publicly available retinal cup and disc segmentation dataset consisting of 750 color fundus images from three different sources: 460 images from MESSIDOR, 195 images from BinRushed, and 95 images from Magrabia. Segmentation masks for the cup and disc contours were manually annotated by six glaucoma specialists following the RIGA benchmarks [28]. During model training, we selected 195 samples from BinRushed and 460 samples from MESSIDOR as the training set as follows [12]. The Magrabia image set containing 95 samples was used as a test set to evaluate the model. We choose the first expert’s annotation as the ground truth for the training set. In order to better evaluate the performance of the method, we construct two test sets, including the mask labeled by the first expert ($\mathbf{Rater1}$) and the mask produced by the majority vote of all experts ($\mathbf{Majority}$ $\mathbf{Vote}$).
@@ III-B Implementation and Evaluation Metrics
We implemented our method and comparative methods in Python using PyTorch and performed the computations on an NVIDIA GeForce RTX 3090 GPU equipped with 8GB of RAM. The backbone network is LinkNet [24] and is trained using the Adam optimizer (betas = (0.9, 0.99)). The initial learning rate was set to 5e-4, divided by 10 for 2000 iterations, with a total of 100 training rounds and an input batch size of 8. The image size of the input network was scaled to 256 × 256 pixels and normalized using the mean and standard deviation of each channel. The model with the best performance on the test set was selected as the final training model. For evaluating the model performance, we give the Dice Similarity Coefficient (DSC) and Intersection over Union (IoU) for each category, excluding the background.
Figure 2: Visualized segmentation results on RIGA of different methods. The ground truth mask and predicted mask are represented by red and green, respectively.
III-C Experiments on RIGA Dataset
Table I presents a comprehensive analysis of the quantitative results obtained from various segmentation methods applied to two distinct test sets (”Majority Vote” and ”Rater1”) within the RIGA dataset. Turning our attention to the performance metrics evaluated, the table showcases results for $IoU_{disc}$, $IoU_{cup}$, $Dice_{disc}$, and $Dice_{cup}$, which serve as key indicators of segmentation accuracy. We observe that LinkNet while delivering competitive performance, falls short of MPNN in terms of all metrics. Co-teaching, TriNet, 2SRnT, MTCL, and PINT also exhibit commendable segmentation results, but they do not surpass the MPNN outcomes. Among the strategies, it’s evident that MPNN emerges as a standout performer. It achieves a remarkable metrics of $IoU_{disc}$ of 85.22%, $IoU_{cup}$ of 78.11%, $Dice_{disc}$ of 91.83%, and $Dice_{cup}$ of 87.25%. Shifting the focus to the ”Rater1” scenario, the trends continue to favor MPNN. Under this scenario, which reflects individual rater annotations, MPNN maintains its performance superiority. With an $IoU_{disc}$ of 83.12%, an $IoU_{cup}$ of 77.34%, a $Dice_{disc}$ of 90.67%, and a $Dice_{cup}$ of 86.77%, MPNN consistently outperforms other methods. In Fig. 2, the visualization results of various methods affirm the efficacy of our proposed approach. These visualizations vividly demonstrate our method’s capacity to precisely segment the targeted regions, aligning seamlessly with the corresponding ground truth annotations. Through these visual representations, we provide qualitative substantiation of our method’s superior performance in contrast to the alternative approaches.
TABLE II: The Ablation Study Of Our Method on RIGA Data. The Impact of Different Numbers of Pseudo-Labels on the Final Result of the Model
@@ III-C Experiments on RIGA Dataset
Methods $K$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
@@
TABLE III: The Ablation Study Of Our Method on RIGA Data. The Impact of Pseudo-Labels Obtained for Different Threshold Training on the Final Results of the Model
@@
Methods $\varphi$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
@@
TABLE IV: The Ablation Study Of Our Method on RIGA Data. Comparison of Using Clean Pixels And Noisy Pixels
Methods $\mathbf{O}_{c}$ $\mathbf{O}_{n}$ Majority Vote
@@
III-D Parameter Ablation Study
In order to determine suitable parameters for the MPNN method on the RIGA dataset, particularly during the pseudo-label generation process, we conducted ablation experiments. The results of our experiments are shown in Table II and Table III. The evaluation is conducted using the ”Majority Vote” scenario, wherein the obtained results for different metrics.
Table II presents the effect of different numbers of pseudo-labels on the segmentation accuracy of MPNN. When considering the metric $IoU_{disc}$, the model’s performance progressively improves from 83.52% at $K=3$ to 84.90% at $K=4$, and further to a notable 85.22% at $K=5$. Similarly, the metric $IoU_{cup}$ shows improvement from 76.94% at $K=3$ to 77.70% at $K=4$, and then to a substantial 78.11% at $K=5$, signifying an improvement of 1.17% from $K=3$ to $K=5$ in the optic disc and cup segmentation IoU. In addition, the Dice coefficients ($Dice_{disc}$ and $Dice_{cup}$) show similar increases. However, when $K=6$, a noteworthy observation emerges: the performance metrics exhibit a slight decline compared to the results obtained at $K=5$. Specifically, for $IoU_{disc}$, the value drops to 85.19%, a decrease of 0.03% from the value at $K=5$. Similarly, for $IoU_{cup}$, the metric decreases to 77.06%, indicating a reduction of 1.05% compared to the value at $K=5$. This observation underlines the importance of striking a balance when introducing pseudo-labels into the training process. While additional pseudo-labels can aid in capturing more label noise, an excessive quantity might lead to the dilution of valuable prior knowledge. This insight encourages further investigation into the optimal amount of pseudo-labels to integrate, considering the trade-off between information enrichment and the risk of introducing noise.
@@ III-D Parameter Ablation Study
Table III presents the impact of multiple pseudo-labels obtained with different training thresholds $\varphi$ on segmentation accuracy in MPNN. The outcomes indicate that the choice of threshold value during pseudo-label generation has a discernible influence on the model’s performance. Notably, for $IoU_{disc}$, we observe a consistent progression as the threshold value increases. The metric shows an improvement from 84.51% at $\varphi=0.91$ to 84.73% at $\varphi=0.92$ and eventually peaks at an impressive 85.22% for $\varphi=0.93$. However, as $\varphi$ increases to 0.94 and 0.95, $IoU_{disc}$ shows a downward trend. When $\varphi$ increases, the resulting clean pixel set may contain inaccurate information, resulting in reduced model performance. Likewise, $IoU_{cup}$ demonstrates a parallel pattern. The performance rises from 76.97% at $\varphi=0.91$ to 76.99% at $\varphi=0.92$ and subsequently reaches its zenith at 78.11% for $\varphi=0.93$. This signifies a prominent advancement of 1.14% in optic disc and cup segmentation IoU between $\varphi=0.91$ and $\varphi=0.93$. In fact, the selection of the threshold is similar to the selection of the number of pseudo-labels, and the trade-off between rich information and the risk of introducing noise needs to be considered. The results affirm the importance of threshold selection during pseudo-label generation, with the highest threshold value of $\varphi=0.93$ yielding the most favorable outcomes in terms of IoU and Dice scores.
III-E Component Ablation Study
To evaluate the effectiveness of each MPNN component on the RIGA dataset, we performed ablation studies using different variants. The results of our ablation experiments are shown in Table IV. The evaluation is conducted using the ”Majority Vote” scenario, wherein the obtained results for different metrics are presented. Table IV presents the comparison of the performance of the MPNN method between different scenes. The symbol $\times$ indicates that the corresponding pixel set is not distinguished when calculating the corresponding loss.
When we input noise pixels and clean pixels without distinguishing between them, the performance of the MPNN is mediocre. As anticipated, when considering only clean pixels or noisy pixels, the MPNN method has demonstrated significant improvement. Furthermore, when both clean pixels and noisy pixels are utilized simultaneously, the MPNN method further demonstrates enhanced results. Specifically, the $IoU_{disc}$ score stands at 85.22%, showcasing an improvement compared to the use of solely clean pixels. Similarly, $IoU_{cup}$ exhibits an increase to 78.11%. The most significant advancements are observed in $Dice_{disc}$, reaching 91.83%, while $Dice_{cup}$ reaches 87.25%. Fig. 3 shows our isolated collection of noisy pixels, mostly at the segmentation edges. In fact, our strategy for selecting multiple pseudo-labels is a form of the ”overcorrection-is-necessary” strategy. The filtered noisy pixels necessarily contain instances that the model can confidently predict. This also explains why introducing additional noisy labels for consistency loss yields better results.
@@ IV Conclusions
In this paper, we introduce a label-denoising method of the Multiple Pseudo-labels Noise-aware Network (MPNN) for precise optic disc and cup segmentation. To distinguish clean pixels and noisy pixels, The proposed MPGGD module uses multiple networks with different initializations to fit the same training set and stops training after reaching a certain threshold to obtain multiple pseudo-labels of the training set. Subsequently, our method can distinguish between noisy and clean pixels by leveraging consensus among multiple pseudo-labels and it proposes distinct learning strategies for noisy and clean pixels based on teacher-student architecture. Through extensive experimentation on the RIGA dataset, our method performs well on the optic disc and cup segmentation tasks. The visualized outcomes further validate the exceptional performance and robust denoising capabilities of our proposed method. Moreover, our approach exhibits superior segmentation efficacy in optic disc and cup segmentation compared to other label-denoising methods, highlighting our capability to precisely capture object boundaries and generate better masks.

### 15 Mar 2024 v2

@@ II-C Final Loss
The total loss function follows the pattern of semi-supervised learning including supervised loss $\mathcal{L}_{cl}$ and unsupervised loss $\mathcal{L}_{no}$. The total loss is expressed as:
$\mathcal{L}_{total}=\beta\mathcal{L}_{cl}+\lambda\mathcal{L}_{no}.$ (7)
Empirically, we set $\beta=1$ and $\lambda$ is a ramp-up trade-off weight commonly scheduled by the time-dependent Gaussian function [18] $\lambda(t)=w_{max}\cdot e^{(-5(1-\frac{t}{t_{max}})^{2})}$, where $w_{max}$ is the maximum weight commonly set as 0.1 [16] and $t_{max}$ is the maximum training iteration. Such a $\lambda$ weight representation avoids being dominated by misleading targets when starting online training.
@@ III-A Dataset and Pre-processing
The RIGA benchmark [19] is a publicly available retinal disc and cup segmentation dataset consisting of 750 color fundus images from three different sources: 460 images from MESSIDOR, 195 images from BinRushed, and 95 images from Magrabia. Segmentation masks for the disc and cup contours were manually annotated by six glaucoma specialists following the RIGA benchmark [19]. We selected 195 samples from BinRushed and 460 samples from MESSIDOR as the training set as follows [6]. The Magrabia image set containing 95 samples was used as a test set to evaluate the model. We choose the first expert’s annotation as the ground truth for the training set. To better evaluate the performance of the method, we construct two test sets, using the masks labeled by the first expert ($\mathbf{Rater1}$) and produced by the majority vote of all experts ($\mathbf{Majority}$ $\mathbf{Vote}$) as ground truth, respectively.
@@ III-B Implementation and Evaluation Metrics
We implemented our method and comparative methods in Python using PyTorch and performed the computations on an NVIDIA GeForce RTX 3090 GPU equipped with 24GB of RAM. We utilized the LinkNet [17] as the backbone and trained it using the Adam optimizer (betas = (0.9, 0.99)). We set the initial learning rate to 5e-4, decreased it by a factor of 10 every 2000 iterations, conducted a total of 100 training rounds, and maintained a batch size of 8. The input network’s image size was scaled to 256 × 256 pixels and normalized using the mean and standard deviation for each channel. To evaluate the model’s performance, we calculated the Intersection over Union (IoU) and Dice for each category, excluding the background.
III-C Experiments on the RIGA Dataset
Table I presents the quantitative results obtained from various segmentation methods applied to two distinct test sets ($\mathbf{Majority}$ $\mathbf{Vote}$ and $\mathbf{Rater1}$). Turning our attention to the performance metrics evaluated, the table showcases results for $IoU_{disc}$, $IoU_{cup}$, $Dice_{disc}$, and $Dice_{cup}$, which serve as key indicators of segmentation accuracy. We observe that LinkNet while delivering competitive performance, falls short of the MPNN in terms of all metrics. Co-teaching, TriNet, 2SRnT, MTCL, and PINT also exhibit commendable results, but they do not surpass the MPNN outcomes. Among the strategies, it’s evident that MPNN emerges as a standout performer. It achieves a remarkable metrics of $IoU_{disc}$ of 85.22%, $IoU_{cup}$ of 78.11%, $Dice_{disc}$ of 91.83%, and $Dice_{cup}$ of 87.25%. Shifting the focus to the scenario of $\mathbf{Rater1}$, the trends continue to favor MPNN. In Fig. 2, the visualization results of various methods affirm the efficacy of the MPNN. These visualizations demonstrate our method’s capacity to precisely segment the targeted regions.
TABLE II: Ablation Study of Our Method on the Majority Vote Test Set. The Impact of Different Numbers of Pseudo-Labels on the Final Results
@@ III-C Experiments on the RIGA Dataset
Method $K$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
@@
TABLE III: Ablation Study of Our Method on the Majority Vote Test Set. The Impact of Pseudo-Labels Obtained for Different Threshold Training on the Final Results
@@
Method $\varphi$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
@@
III-D Parameter Ablation Study
To determine suitable parameters for the MPNN on the RIGA dataset, particularly during the pseudo-labels generation process, we conducted ablation experiments in the scenario of $\mathbf{Majority}$ $\mathbf{Vote}$.
Table II presents the effect of different numbers of pseudo-labels on the segmentation accuracy of the MPNN. When considering the metric $IoU_{disc}$, the model’s performance progressively improves from 83.52% at $K=3$ to 84.90% at $K=4$, and further to a notable 85.22% at $K=5$. Similarly, the metric $IoU_{cup}$ shows improvement from 76.94% at $K=3$ to 77.70% at $K=4$, and then to a substantial 78.11% at $K=5$. In addition, the Dice coefficients ($Dice_{disc}$ and $Dice_{cup}$) show similar increases. However, when $K=6$, a noteworthy observation emerges: the performance metrics exhibit a slight decline compared to the results obtained at $K=5$. This observation underlines the importance of striking a balance when introducing pseudo-labels into the training process. While additional pseudo-labels can aid in capturing more label noise, an excessive quantity might lead to the dilution of valuable prior knowledge. This insight encourages further investigation into the optimal amount of pseudo-labels to integrate, considering the trade-off between information enrichment and the risk of introducing noise.
Table III presents the impact of multiple pseudo-labels obtained with different training thresholds $\varphi$ on the segmentation accuracy of MPNN. Notably, for $IoU_{disc}$, we observe a consistent progression as the threshold value increases. The metric shows an improvement from 84.51% at $\varphi=0.91$ to 84.73% at $\varphi=0.92$ and eventually peaks at an impressive 85.22% for $\varphi=0.93$. However, as $\varphi$ increases to 0.94 and 0.95, $IoU_{disc}$ shows a downward trend. When $\varphi$ increases, the resulting clean pixel set may contain inaccurate information, resulting in reduced model performance. The results affirm the importance of threshold selection during pseudo-label generation, with the highest threshold value of $\varphi=0.93$ yielding the most favorable outcomes in terms of IoU and Dice scores.
TABLE IV: Ablation Study of Our Method on the Majority Vote Test Set. Comparison Results of Using Clean Pixels And Noisy Pixels
Method $\mathbf{O}_{cl}$ $\mathbf{O}_{no}$ Majority Vote
@@ III-D Parameter Ablation Study
III-E Component Ablation Study
To evaluate the effectiveness of each MPNN component, we performed ablation studies using different variants. The results of our ablation experiments are shown in Table IV. The evaluation is conducted in the scenario of $\mathbf{Majority}$ $\mathbf{Vote}$. The symbol $\times$ indicates that the corresponding pixel set is not distinguished when calculating the corresponding loss.
When we input noisy pixels and clean pixels without distinguishing between them, the performance of the MPNN is mediocre. As anticipated, when considering only clean pixels or noisy pixels, the MPNN method has demonstrated significant improvement. Furthermore, when both clean pixels and noisy pixels are utilized simultaneously, the MPNN method further demonstrates enhanced results. Fig. 3 shows our isolated collection of noisy pixels, mostly at the segmentation edges. Our strategy for selecting multiple pseudo-labels is a form of the “overcorrection-is-necessary” strategy. The filtered noisy pixels necessarily contain instances that the model can confidently predict. This also explains why introducing additional noisy labels for consistency loss yields better results.
@@ IV Conclusions
In this paper, we introduce a label-denoising method of the Multiple Pseudo-labels Noise-aware Network (MPNN) for precise optic disc and cup segmentation. The proposed MPGGD module uses multiple networks with different initializations to fit the same training set and stops training after reaching a certain threshold to obtain multiple pseudo-labels of the training set. Subsequently, our method can distinguish between noisy and clean pixels by leveraging consensus among multiple pseudo-labels and it proposes distinct learning strategies for noisy and clean pixels based on teacher-student architecture. Through extensive experimentation on the RIGA dataset, our method performs well on the optic disc and cup segmentation tasks. The visualized outcomes further validate the exceptional performance and robust denoising capabilities of our proposed method. However, there are significant differences among various medical image segmentation tasks with ambiguous boundaries, and such variations can noticeably impact algorithm performance. Therefore, the successful extension of the proposed MPNN to other domains of medical image segmentation with ambiguous boundaries is a direction worthy of in-depth exploration.
V Acknowledge
This research was supported by the Zhejiang Provincial Natural Science Foundation of China under Grant No. LDT23F01015F01 and No. LY21F020004, Natural Science Foundation of Jiangsu Province under Grant No. BK20220266, and National Natural Science Foundation of China No. 61572243 and Construction Project of Lishui University Discipline (Zhejiang Province First Class Discipline, Discipline name: intelligent science and technology) No. XK0430403005.